5. 6. At a local high school 5000 juniors and seniors recently took an aptitude test. The results of the exam were normally d 5-8 show work
Answer #1

Total Sooo students e=so mean M = 450 X - Normal distribution z = X-2 la) over uas P(x> 425) = P(2> 425-1 - Plz> 425-459 = P(